# GMIMS-DRAGONS: A Faraday Depth Survey of the Northern Sky Covering 350 to 1030 MHz

### v1

@@ Abstract
Polarized synchrotron emission at meter to centimeter wavelengths provides an effective tracer of the Galactic magnetic field. Calculating Faraday depth, the most useful parameter for mapping the line-of-sight magnetic field, requires observations covering wide frequency bands with many channels. As part of the Global Magneto-Ionic Medium Survey we have observed polarized emission spanning 350–1030 MHz over the Northern sky, in the declination range ${-20\arcdeg}\leq{\delta}\leq{90\arcdeg}$. We used the 15 m telescope at the Dominion Radio Astrophysical Observatory (DRAO), equipped to receive orthogonal circular polarizations, with the Onsala Space Observatory band-1 feed developed for the Square Kilometre Array. Angular resolution varies across the band from $1.3\arcdeg$ to $3.6\arcdeg$. A digital spectrometer provided 42 kHz frequency resolution. Data were taken with the telescope moving rapidly in azimuth and are absolutely calibrated in intensity. Approximately 20% of the band was obscured by radio-frequency interference. Resolution in Faraday depth is $\sim 6$ rad m-2, and features as wide as $\sim 38$ rad m-2 are represented. The median sensitivity of the Faraday depth cube is 11 mK. Approximately 55% of sight-lines in this survey show Faraday complexity. This dataset, called “DRAO GMIMS of the Northern Sky” (DRAGONS), is the first to probe Faraday depth of the Northern sky in its frequency range and will support many scientific investigations. The data will be used to calibrate surveys with higher angular resolution, particularly Galactic foreground maps from the Canadian Hydrogen Intensity Mapping Experiment, and to provide information on large structures for aperture-synthesis telescopes, particularly the DRAO Synthesis Telescope. The data are available through the Canadian Astronomy Data Centre.
@@
facilities: DRAO:DRAO-15software: astropy (Astropy Collaboration et al., 2022); matplotlib (Hunter, 2007); numpy; RM-tools (Purcell et al., 2020); healpy (Zonca et al., 2019; Górski et al., 2005); ALBUS.
1 Introduction
Magnetic fields play a significant role in the physics of the Galaxy, contributing to the formation of clouds and stars (e.g., Federrath & Klessen, 2012; Tahani et al., 2022a, b), and influencing Galactic evolution (Kim et al., 1996). The magnetic field shares the energy of the interstellar medium (ISM) in approximately equal parts (on sufficiently large scales) with the energetic charged particles and the motion of the gas (Ferrière, 2001). The interplay between magnetic fields, cosmic rays, gas, dust and stars is responsible for the structures appearing in the ISM on a wide range of spatial scales. Mapping the magnetic field across the sky, and along the line of sight (LOS), will provide fundamental information on the Galaxy and its future development. This three dimensional (3D) mapping is challenging, but important, and high quality, all-sky data are necessary for constraining global, large-scale models (e.g., Unger & Farrar, 2024). Recent studies have shown that mapping 3D magnetic field vectors of Galactic objects such as molecular clouds reveals their formation and evolution history (e.g. Tahani et al., 2022a, b).
@@ 1 Introduction
where $n_{e}$ is the electron number density and $B_{||}$ is the LOS component of the magnetic field, which is positive when directed toward the observer. We follow the convention in Equation 17 of Ferrière et al. (2021), with the integral calculated along the LOS path length increments, $dr$, from the observer located at $r=0$ to the polarized emission source located at a distance $r=d$.
In the simplest case of an emission source located at a discrete distance experiencing Faraday rotation through a magnetized, ionized medium between the source and observer, the LOS is characterized by a single Faraday depth, known as a rotation measure (RM), which can be calculated from a linear relationship between the measured polarization angle and the square of the wavelength, $\lambda$. However, in the case of the diffuse polarized synchrotron emission in the ISM, emission and Faraday rotation can be mixed, with sources located over a range of different distances, $d$, all undergoing the Faraday rotation caused by varying pathlengths through the intervening medium. The resulting effects on the observed, complex polarization can, in part, be disentangled by applying ‘Faraday synthesis’ (also known as ‘RM synthesis’; Burn, 1966; Brentjens & de Bruyn, 2005), which recovers a spectrum of Faraday depths describing the LOS instead of a single value. The complex Faraday depth spectrum, $\tilde{P}(\phi)$, can be calculated as
@@
where $\tilde{P}(\lambda^{2})$ is the measured complex polarization across the available observing frequencies (wavelengths) and $W(\lambda^{2})$ is a weighting function representing the finite $\lambda^{2}$ coverage. The range of scales in Faraday depth space that can be recovered using this method depends on the observing bandwidth and the width and sampling of the frequency channels. Broad frequency coverage (octave bandwidth) and narrow channel widths ($\sim 1$ MHz or less) are best suited to maximizing the information that can be gained from measuring Faraday rotation.
Faraday rotation is a tracer of both the magnetic field and the ionized gas density. Recent studies have shown that Faraday rotation is sensitive to relatively low electron densities, meaning that it is able to probe not only the warm ionized medium, but also the warm partially-ionized medium (Heiles & Haverkorn, 2012) and even the low ionization of the warm neutral medium (Foster et al., 2013; Van Eck et al., 2017; Bracco et al., 2022), particularly at low frequencies (<1<1 GHz).
The Global Magneto-Ionic Medium Survey (GMIMS) is paving the way for full-sky, broadband, single-antenna polarization maps that are ideal for the application of the Faraday synthesis technique, enabling the study of complex effects of mixed synchrotron emission and Faraday rotation in the ISM. The GMIMS initiative is loosely comprised of six component surveys, divided into three bands in each of the northern and southern hemispheres. Three of these surveys have been published: (i) GMIMS-Low-Band South (GMIMS-LBS), covering 300–480 MHz (Wolleben et al., 2019), (ii) the Southern Twenty Centimeter All-sky Polarization Survey (STAPS), contributing GMIMS-High-Band South (GMIMS-HBS) and covering 1324–1770 MHz (Sun et al., 2025), both using Murriyang, CSIRO’s Parkes 64 m radio telescope, and (iii) GMIMS-High-Band-North (GMIMS-HBN), covering 1280–1750 MHz (Wolleben et al., 2021) using the Dominion Radio Astrophysical Observatory (DRAO) 26 m John A. Galt telescope. Of the remaining planned surveys, observations are close to completion for the POSSUM EMU GMIMS All-Stokes UWL Survey (PEGASUS; E. Carretti et al., 2026, in preparation), covering 704–1440 MHz and contributing GMIMS-Mid-Band-South, using the Murriyang telescope at Parkes.
Interferometric data from several other instruments and surveys are complementary to the GMIMS single-antenna datasets. A polarization study using data from the Canadian Hydrogen Intensity Mapping Experiment (CHIME; CHIME Collaboration et al., 2022) has been published (Mohammed et al., 2024), demonstrating great promise for incorporating CHIME maps into GMIMS as a high-resolution component in the northern hemisphere, covering 400–800 MHz. Separately, the LOFAR Two-meter Sky Survey (LoTSS-DR2) has demonstrated a low-frequency (120–167 MHz), interferometric approach to diffuse emission Faraday synthesis maps over a significant portion of the northern sky (Erceg et al., 2022, 2024). The Polarisation Survey of the Universe’s Magnetism (POSSUM) is also producing high-resolution diffuse polarization maps (Gaensler et al., 2025), for which PEGASUS will ultimately provide the short baselines.
@@
Here we present the DRAO GMIMS of the Northern Sky (DRAGONS) survey, covering 350–1030 MHz, which contributes GMIMS-Low-Band-North (GMIMS-LBN). The frequency range corresponds to a $\lambda^{2}$ range of 0.085 m2 to 0.73 m2, the widest $\lambda^{2}$ coverage of all GMIMS surveys, allowing for unprecedented sensitivity to Faraday complexity. The overlap in sky coverage ($20\arcdeg\leq\delta\leq 20\arcdeg$) and frequency (350–480 MHz) with GMIMS-LBS allows for useful comparison and data validation between the two low-frequency GMIMS surveys. The complete sky-coverage overlap with CHIME data, and 400–800 MHz frequency overlap, along with complementary sensitivity to different ranges of angular scales, will allow DRAGONS to be used to calibrate CHIME polarization maps and provide information on missing short baselines.
@@
This paper is organized as follows. In Section 2 we describe the DRAO 15 m telescope, its signal chain, and the science commissioning steps. In Section 3 we outline the observing strategy and describe the data collected for the survey and calibration. In Section 4 we describe the data processing pipeline and map-making steps. In Section 5 we present Stokes $I$, $Q$ and $U$ maps, and a data quality assessment with comparisons to previously published data sets. In Section 6 we describe the Faraday synthesis procedures and present the resulting Faraday depth cubes. In Section 7 we discuss some preliminary science results, with comparisons to other GMIMS data products. We conclude in Section 8 with a summary of DRAGONS science papers currently underway, and a look toward future work with the broadband capabilities of this survey.
2 Telescope and Receiver
@@
The telescope we use was originally proposed as a Square Kilometre Array (SKA) design. Although not selected for the SKA, the prototype instrument, installed at the South-West corner of the DRAO site, has proven to be useful as a single-antenna telescope with its rapid survey capability. We describe the components comprising the telescope below, including the unique reflector configuration, the feed and receiver, and the signal chain. We also describe the characterization of the telescope completed during the science commissioning phase preceding this survey.
2.1 The 15 m Telescope
@@
We conducted survey observations with the ${15}\times{18}$ m offset Gregorian reflector at the Dominion Radio Astrophysical Observatory (DRAO), pictured in Figure 1. The telescope was designed to obtain an unblocked aperture of diameter 15 m (Lacy et al., 2012; Knee et al., 2016). We refer to it as the DRAO-15. Both reflecting surfaces are built from carbon fiber, with surface precision ${\sim}250~{\mu}{\rm{m}}$. The antenna was equipped to receive left- and right-hand circular polarization (denoted here by $L$ and $R$ respectively). Offset reflectors have poor polarization performance, with a position offset between $L$ and $R$ beams (Chu & Turrin, 1973), but the problem is overcome in this telescope by Baker & Imbriale (2014); Baker (2020) through shaping of the reflector surfaces following the technique described by Mizugutch et al. (1976, hereafter ‘Mizigutch technique’). Further shaping of the reflector surfaces is used to enhance aperture efficiency.
The reflector was fed with a quad-ridged, flared horn (Flygare et al., 2018) with an integrated calibration noise coupler (Flygare et al., 2017), the prototype feed developed by Chalmers University’s Onsala Space Observatory (OSO) for the SKA-mid Band 1. The feed accepts two orthogonal linearly polarized signals. To measure the linearly polarized Galactic emission, circular polarization is preferred over linear because Stokes parameters $Q$ and $U$ can then both be derived using correlation techniques, as described in McConnell et al. (2006) and Robishaw & Heiles (2021). This method is critically important for single-antenna systems, as it avoids the inherent instability of deriving Stokes $Q$ from the difference of two large, noisy total-power signals. By converting to a circular basis, both Stokes $Q$ and $U$ are instead derived from the more stable cross-correlation product, which is largely immune to uncorrelated receiver noise and gain fluctuations. To achieve this, the linearly polarized outputs from the feed were connected to a transmission-line hybrid coupler [^1] R&D Microwaves model HD-A01 before amplification so that the telescope outputs became $L$ and $R$. Losses in transmission lines, in connectors, and in the hybrid amount to ${\sim}12$ K at 350 MHz and ${\sim}35$ K at 1030 MHz.
Figure 2 shows the measured half-power beamwidth (HPBW) across the band. The HPBW varies from $3.6^{\circ}$ at 350 MHz to $1.3^{\circ}$ at 1030 MHz (all frequency channels maps are convolved to a common $3.6\arcdeg$ resolution prior to Faraday synthesis; Section 6.1). Figure 3 shows the positions of Cyg A derived from raster scans (described in Section 3.2.2), averaged over seven scans. The top two panels show the difference between the expected and measured positions. There is some variation of beam position with frequency, but the offsets are typically less than 5% of the beamwidth. The bottom two panels of Figure 3 show the offset between the RR* and LL* beams. These measurements show the extent to which the offset reflector, corrected by the Mizigutch technique, provides coincident beams in the two hands of circular polarization. This small offset has a negligible effect on our data.
@@
Figure 2: The measured HPBW of the telescope as a function of frequency, determined from the FWHM of Gaussians fitted to raster scans of compact sources (see Section 3.2.2). Gaps in the measured beamwidth data are channels with too much RFI for a successful Gaussian fit. A boxcar median smoothing kernel of width 49 channels ($\sim 4$ MHz) was applied to the measured FWHM values.
@@
Figure 3: Frequency dependence of pointing measurements using the average of seven Cyg A raster scans. (a) Right ascension offset, (b) declination offset, (c) difference in right ascension measured in LL* and RR*, and (d) difference in declination measured in LL* and RR*. The consistently positive offsets in (a) and (b) are the residual, systematic errors in the pointing accuracy following the correction described in Section 2.5. The ‘persistent RFI’ mask described in Section 4.1 was applied to these data to eliminate most RFI channels. Remaining outliers are due to poor-quality fits to the raster scans in channels contaminated by intermittent RFI.
2.2 The 350–1030 MHz Receiver

### v2

# GMIMS-DRAGONS: A Faraday Depth Survey of the Northern Sky Covering 350–1030 MHz

@@ Abstract
Polarized synchrotron emission at meter to centimeter wavelengths provides an effective tracer of the Galactic magnetic field. Calculating Faraday depth, the most useful parameter for mapping the line-of-sight magnetic field, requires observations covering wide frequency bands with many channels. As part of the Global Magneto-Ionic Medium Survey (GMIMS), we have observed polarized emission spanning 350–1030 MHz over the northern sky, in the declination range ${-20\arcdeg}\leq{\delta}\leq{90\arcdeg}$. We used the 15 m telescope at the Dominion Radio Astrophysical Observatory (DRAO), equipped to receive orthogonal circular polarizations, with the Onsala Space Observatory band 1 feed developed for the SKA Project. Angular resolution varies across the band from $1.3\arcdeg$ to $3.6\arcdeg$. A digital spectrometer provided 42 kHz frequency resolution. Data were taken with the telescope moving rapidly in azimuth and are absolutely calibrated in intensity. Approximately 25% of the data were lost due to radio-frequency interference. The resolution in Faraday depth is $\sim 6$ rad m-2, and features as wide as $\sim 38$ rad m-2 are represented. The median sensitivity of the Faraday depth cube is 11 mK. Approximately 55% of sight-lines in this survey show Faraday complexity. This dataset, called “DRAO GMIMS of the Northern Sky” (DRAGONS), is the first to probe Faraday depth of the northern sky in its frequency range and will support many scientific investigations. The data will be used to calibrate surveys with higher angular resolution, particularly Galactic foreground maps from the Canadian Hydrogen Intensity Mapping Experiment, and to provide information on large structures for aperture-synthesis telescopes, particularly the DRAO Synthesis Telescope. The data are available through the Canadian Astronomy Data Centre.
@@
facilities: DRAO:15msoftware: astropy (Astropy Collaboration et al., 2022); matplotlib (Hunter, 2007); numpy; RM-tools (Purcell et al., 2020); healpy (Zonca et al., 2019; Górski et al., 2005); ALBUS.
I Introduction
Magnetic fields play a significant role in the physics of the Galaxy, contributing to the formation of clouds and stars (e.g., Federrath and Klessen, 2012; Tahani et al., 2022b, a), and influencing Galactic evolution (Kim et al., 1996). The magnetic field shares the energy of the interstellar medium (ISM) in approximately equal parts (on sufficiently large scales) with the energetic charged particles and the motion of the gas (Ferrière, 2001). The interplay between magnetic fields, cosmic rays, gas, dust, and stars is responsible for the structures appearing in the ISM on a wide range of spatial scales. Mapping the magnetic field across the sky, and along the line of sight (LOS), will provide fundamental information on the Galaxy and its future development. This 3D mapping is challenging, but important, and high quality, all-sky data are necessary for constraining global, large-scale models (e.g., Unger and Farrar, 2024). Recent studies have shown that mapping 3D magnetic field vectors of Galactic objects such as molecular clouds reveals their formation and evolution history (e.g., Tahani et al., 2022b, a).
@@ I Introduction
where $n_{e}$ is the electron number density, and $B_{||}$ is the LOS component of the magnetic field, which is positive when directed toward the observer. We follow the convention in Equation 17 of Ferrière et al. (2021), with the integral calculated along the LOS path length increments, $dr$, from the observer located at $r=0$ to the polarized emission source located at a distance $r=d$.
In the simplest case of an emission source located at a discrete distance experiencing Faraday rotation through a magnetized, ionized medium between the source and observer, the LOS is characterized by a single Faraday depth, known as a rotation measure (RM), which can be calculated from a linear relationship between the measured polarization angle and the square of the wavelength, $\lambda$. However, in the case of the diffuse polarized synchrotron emission in the interstellar medium (ISM), emission and Faraday rotation can be mixed, with sources located over a range of different distances, $d$, all undergoing the Faraday rotation caused by varying pathlengths through the intervening medium. The resulting effects on the observed, complex polarization can, in part, be disentangled by applying ‘Faraday synthesis’ (also known as ‘RM synthesis’; Burn, 1966; Brentjens and de Bruyn, 2005), which recovers a spectrum of Faraday depths describing the LOS instead of a single value. The complex Faraday depth spectrum, $\tilde{P}(\phi)$, can be calculated as
@@
where $\tilde{P}(\lambda^{2})$ is the measured complex polarization across the available observing frequencies (wavelengths), and $W(\lambda^{2})$ is a weighting function representing the finite $\lambda^{2}$ coverage. The range of scales in Faraday depth space that can be recovered using this method depends on the observing bandwidth and the width and sampling of the frequency channels. Broad frequency coverage (octave bandwidth) and narrow channel widths ($\sim 1$ MHz or less at frequencies below 2 GHz) are best suited to maximizing the information that can be gained from measuring Faraday rotation.
Faraday rotation is a tracer of both the magnetic field and the ionized gas density. Recent studies have shown that Faraday rotation is sensitive to relatively low electron densities, meaning that it is able to probe not only the warm ionized medium, but also the warm partially-ionized medium (Heiles and Haverkorn, 2012) and even the low ionization of the warm neutral medium (Foster et al., 2013; Van Eck et al., 2017; Bracco et al., 2022), particularly at low frequencies (<1<1 GHz).
The Global Magneto-Ionic Medium Survey (GMIMS) is paving the way for full-sky, broadband, single-antenna polarization maps that are ideal for the application of the Faraday synthesis technique, enabling the study of complex effects of mixed synchrotron emission and Faraday rotation in the ISM. The GMIMS initiative is loosely comprised of six component surveys, divided into three bands in each of the northern and southern hemispheres. Three of these surveys have been published: (i) GMIMS-Low-Band-South (GMIMS-LBS), covering 300–480 MHz (Wolleben et al., 2019), (ii) the Southern Twenty Centimeter All-sky Polarization Survey (STAPS), contributing GMIMS-High-Band-South (GMIMS-HBS) and covering 1324–1770 MHz (Sun et al., 2025), both using Murriyang, CSIRO’s Parkes 64 m radio telescope, and (iii) GMIMS-High-Band-North (GMIMS-HBN), covering 1280–1750 MHz (Wolleben et al., 2021) using the Dominion Radio Astrophysical Observatory (DRAO) 26 m John A. Galt telescope. Of the remaining planned surveys, observations are close to completion for the POSSUM (Polarisation Sky Survey of the Universe’s Magnetism) EMU (Evolutionary Map of the Universe) GMIMS All-Stokes UWL (Ultra Wideband Low) Survey (PEGASUS; E. Carretti et al. 2026, in preparation), covering 704–1440 MHz and contributing GMIMS-Mid-Band-South, using the Murriyang telescope at Parkes.
@@
Interferometric data from several other instruments and surveys are complementary to the GMIMS single-antenna datasets. A polarization study using data from the Canadian Hydrogen Intensity Mapping Experiment (CHIME; CHIME Collaboration et al., 2022) has been published (Mohammed et al., 2024), demonstrating great promise for incorporating CHIME maps into GMIMS as a high-resolution component in the northern hemisphere, covering 400–800 MHz. Separately, the LOFAR Two-meter Sky Survey (LoTSS-DR2) has demonstrated a low-frequency (120–167 MHz), interferometric approach to diffuse emission Faraday synthesis maps over a significant portion of the northern sky (Erceg et al., 2022, 2024). POSSUM is also producing high-resolution diffuse emission polarization maps (Gaensler et al., 2025), for which PEGASUS will ultimately provide the short baselines.
@@
Here we present the DRAO GMIMS of the Northern Sky (DRAGONS) survey, covering 350–1030 MHz, which contributes GMIMS-Low-Band-North. The frequency range corresponds to a $\lambda^{2}$ range of 0.085 m2 to 0.73 m2, the widest $\lambda^{2}$ coverage of all GMIMS surveys, allowing for unprecedented sensitivity to Faraday complexity. The overlap in sky coverage ($20\arcdeg\leq\delta\leq 20\arcdeg$) and frequency (350–480 MHz) with GMIMS-LBS allows for a useful comparison and data validation between the two low-frequency GMIMS surveys. The complete sky-coverage overlap with CHIME data, and 400–800 MHz frequency overlap, along with complementary sensitivity to different ranges of angular scales, will allow DRAGONS to be used to calibrate CHIME polarization maps and provide information on missing short baselines.
@@
This paper is organized as follows. In Section II we describe the DRAO 15 m telescope, its signal chain, and the science commissioning steps. In Section III we outline the observing strategy and describe the data collected for the survey and calibration. In Section IV we describe the data processing pipeline and map-making steps. In Section V we present Stokes $I$, $Q$, and $U$ maps, and a data quality assessment with comparisons to previously published data sets. In Section VI we describe the Faraday synthesis procedures and present the resulting Faraday depth cubes. In Section VII we discuss some preliminary science results, with comparisons to other GMIMS data products. We conclude in Section VIII with a summary of DRAGONS science papers currently underway, and a look toward future work with the broadband capabilities of this survey.
II Telescope and Receiver
@@
The telescope we use was originally proposed as an SKA Project design. Although not selected for the SKA, the prototype instrument, installed at the south-west corner of the DRAO site, has proven to be useful as a single-antenna telescope with its rapid survey capability. We describe the components comprising the telescope below, including the unique reflector configuration, the feed and receiver, and the signal chain. We also describe the characterization of the telescope completed during the science commissioning phase preceding this survey.
II.1 The 15 m Telescope
@@
We conducted survey observations with the ${15}\times{18}$ m offset Gregorian reflector at the DRAO, pictured in Figure 1. The telescope was designed to obtain an unblocked aperture of 15 m diameter (Lacy et al., 2012; Knee et al., 2016). We refer to it as the DRAO-15. Both reflecting surfaces are built from carbon fiber, with a surface rms ${\sim}250~{\mu}{\rm{m}}$. We equipped the telescope to receive left- and right-hand circular polarization (denoted here by $L$ and $R$ respectively). Offset reflectors have poor polarization performance, with a position offset between $L$ and $R$ beams (Chu and Turrin, 1973), but the problem is overcome in this telescope by Baker and Imbriale (2014); Baker (2020) through shaping of the reflector surfaces following the technique described by Mizugutch et al. (1976, hereafter ‘Mizigutch technique’). Further shaping of the reflector surfaces is used to enhance aperture efficiency.
@@
Figure 2: The measured HPBW of the telescope as a function of frequency, determined from the FWHM of Gaussians fitted to raster scans of compact sources (see Section III.2.2). Gaps in the measured beamwidth data are channels with too much RFI for a successful Gaussian fit. A boxcar median smoothing kernel of width 49 channels ($\sim 4$ MHz) was applied to the measured FWHM values.
@@
Figure 3: Frequency dependence of pointing measurements using the average of seven Cyg A raster scans. (a) Right ascension offset, (b) declination offset, (c) difference in right ascension measured in LL* and RR*, and (d) difference in declination measured in LL* and RR*. The consistently positive offsets in (a) and (b) are the residual, systematic errors in the pointing accuracy following the correction described in Section II.5. The ‘persistent RFI’ mask described in Section IV.1 was applied to these data to eliminate most RFI channels. Remaining outliers are due to poor-quality fits to the raster scans in channels contaminated by intermittent RFI.
The reflector was fed with a quad-ridged, flared horn (Flygare et al., 2018) with an integrated calibration noise coupler (Flygare et al., 2017), the prototype feed developed by Chalmers University’s Onsala Space Observatory (OSO) for the SKA-mid Band 1. The feed accepts two orthogonal linearly polarized signals. To measure the linearly polarized Galactic emission, circular polarization is preferred over linear because Stokes parameters $Q$ and $U$ can then both be derived using correlation techniques, as described in McConnell et al. (2006) and Robishaw and Heiles (2021). This method is critically important for single-antenna systems, as it avoids the inherent instability of deriving Stokes $Q$ from the difference of two large, noisy total-power signals. By converting to a circular basis, both Stokes $Q$ and $U$ are instead derived from the more stable cross-correlation product, which is largely immune to uncorrelated receiver noise and gain fluctuations. To achieve this, the linearly polarized outputs from the feed were connected to a transmission-line hybrid coupler [^1] R&D Microwaves model HD-A01 before amplification so that the telescope outputs became $L$ and $R$. Losses in transmission lines, in connectors, and in the hybrid amount to ${\sim}12$ K at 350 MHz and ${\sim}35$ K at 1030 MHz.
Figure 2 shows the measured half-power beamwidth (HPBW) across the band. The HPBW varies from $3.6^{\circ}$ at 350 MHz to $1.3^{\circ}$ at 1030 MHz (all frequency channel maps are convolved to a common $3.6\arcdeg$ resolution prior to Faraday synthesis; Section VI.1). Figure 3 shows the positions of Cyg A derived from raster scans (described in Section III.2.2), averaged over seven scans. The top two panels show the difference between the expected and measured positions. There is some variation of beam position with frequency, but the offsets are typically less than 5% of the beamwidth. The bottom two panels of Figure 3 show the offset between the RR* and LL* beams. These measurements show the extent to which the offset reflector, corrected by the Mizigutch technique, provides coincident beams in the two hands of circular polarization. This small offset has a negligible effect on our data.
II.2 The 350–1030 MHz Receiver